# Exceptional topology in Non-Hermitian Twisted Bilayer Graphene

### v2

@@ IV Exceptional cone
To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. 1 at the first honeycomb shell. The Hamiltonian is written as
@@
where $h_{tj}$ for $j=1,2,3$ are on the top layer and $h_{b}$ on the bottom layer. Since $h_{b}$ is non-Hermitian, the Hamiltonian $H_{tri}$ has biorthogonal eigenvectors. In this system, the left and right eigenvectors are four two-component spinors
@@
Using perturbation theory one can derived the effective Hamiltonian fin the space of $\psi_{0}$ The Hamiltonian $H_{tri}$ can be divide into a momentum-independent part $\mathcal{H}^{(0)}$ and a momentum-dependent part $\mathcal{H}^{(1)}_{k}$. The $\psi_{jR}$ can be expressed by
@@
Since $h_{j}$ is independent of non-Hermicity, we can expect $T_{j}h_{j}^{-1}T_{j}=0$, which is similar to the Hermitian model. Thus, the $\psi_{0}$ spinor in the zero-energy eigenstates satisfies
@@
This form seems to be the same as that of the Hermitian TBG. However, $h_{0}$ is in a non-Hermitian form, the $\psi_{0R}$ is different from $\psi_{0L}$. We can check that the biorthorgonal wave functions are normalized as
@@
The effective Hamiltonian matrix to leading order in $k$ is
$\langle\Psi^{(i)}|\mathcal{H}^{(1)}_{\textbf{k}}|\Psi^{(j)}\rangle=\psi_{0}^{(i){\dagger}}(-v_{F}^{*}\rm{\sigma}\cdot\mathbf{k}+i\lambda_{V}^{*}\sigma_{z})\psi^{(j)}_{0}$ (10)
@@
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in AB region $\alpha_{1}$. To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases for $\alpha_{0}=0$ and $\alpha_{0}\neq 0$. When $\alpha_{0}=0$, $\lambda_{V}^{*}$ increases with increasing interlayer coupling in AB region $\alpha_{1}$. Thus, the ratio $\lambda_{V}^{*}/v_{F}^{*}$ becomes larger as $\alpha_{1}$ increases. When $\alpha_{0}\neq 0$, $\lambda_{V}^{*}$ increases with increasing interlayer coupling in AA region $\alpha_{0}$ and the non-Hermitian effect is more complicated. We can expect that the exceptional magic angle $\alpha_{1c}$ is not only determined by $\lambda_{V}$ but also $\alpha_{0}$.
V Berry Curvature and Chern number
@@
So far, we have shown the existence of exceptional geometries in the non-Hermitian TBG. In this section, we will explore their topological properties numerically.
@@
Here, $|\psi^{\beta^{\prime}}(\mathbf{k})\rangle$ and $\langle\psi^{\beta}(\mathbf{k})|$ can be the left or right eigenvectors. Since previous study has proven that the total topological charge is the same for four possible choices of eigenvectors [47], we will choose $\beta=L$ and $\beta^{\prime}=R$ in the following calculation.
Figure 4: (color online) Berry curvature for the $\sigma_{z}=1$ band in the moiré Brllouin zone for $u_{0}=0$, $u_{1}=0.005$, $\theta=1$ and $\lambda_{V}=0.08>\lambda_{Vc}$. A small in-plane electric field is imposed to open the gap.
@@
$\Omega(\mathbf{k})^{\beta,\beta^{\prime}}=\lim_{q\rightarrow 0}\frac{1}{4q^{2}}\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}|\psi^{\beta}_{\mathbf{k}-q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}+q\hat{y}}|\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}+q\hat{x}-q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}-q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}\rangle$ (15)
@@
Figure 4 shows the Berry curvature in mBZ. We can see that Berry curvatures have sharp peaks around $\Gamma_{M}-K_{M}$ high symmetric lines, which are close to the boundary of exceptional rings. This can be well understood since Berry curvature is the phase of an overlap tracing a small loop in mBZ. On the boundary of exceptional rings, the wave functions vary rapidly and will give a large Berry curvature.
@@
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. For small $\lambda_{V}$, we can find that the total Berry charge in one mBZ is 1. This can be easy to understand from two facts. One is that each of the three Dirac points corresponds to Berry curvature monopoles of charge 1 in the absence of non-Hermiticity. The other is that the topological charge is preserved on the exceptional contour that forms from the original Dirac point. Thus, we can expect that each of the three exceptional contours contributes one-thirds of the Berry charge to one mBZ.
@@
Most interestingly, even when $\lambda_{V}$ exceeds $\lambda_{Vc}$, the total Berry charge remains one. This is opposite to the merging of exceptional contours with opposite topological charge [36], in which the topological charge is dissipated and gives a single, uncharged exceptional contour. This indicates a breakdown of Nielsen-Ninomiya theorem and distinguishes the valley topological features of non-Hermitian TBG from a single-layer graphene.
At much larger $\lambda_{V}$ (e.g., $\lambda_{V}>0.85$ for $\theta=1$), the Berry charge can be a quantized value larger than 1 due to the band crossings with the remote bands.
VI Discussion and conclusion
@@
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings centered at $K_{M}$ ($K^{\prime}_{M}$) corner in the moiré Brillouin zone in the $K^{\prime}$ ($K$) valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This flat band is characterized with a robust Chern number $\mathcal{C}=1$ in moiré Brillouin zone. This is different from single-layer graphene in which Dirac/Weyl exceptional rings carrying opposite topological charges are coupled and annihilated.
@@
We have introduced a new design concept for twisted moiré systems, in which a certain “exceptional Dirac magic” angle is achieved. Non-Hermitian TBG can be realized in cold atom and metamaterial system. On one hand, due to the great advantage of manipulating atom-atom interaction and light-matter coupling, ultracold atom experiments provide an unprecedented opportunity. Remarkably, atomic Bose-Einstein condensate has been recently realized in twisted bilayer optical lattices  [48]. Also, there are proposals on the photonic analog [49, 50, 51, 52, 53] and phononic analog [54, 55] of twisted bilayer graphene. In particular, there is demonstration of optical bilayer photonic crystal devices in the microwave range [56] and in the optical frequency range [57]On the other hand, on-site dissipations are well-controlled non-Hermitian terms that can be realized experimentally in many systems [32, 33, 34]. In particular, the Weyl exceptional ring has been realized in photonic experiment [58, 59].
The highlight of this work is the discovery of non-dissipation exceptional topology. The interplay of topology, non-Hermitian, and moiré physics can open the door to new phenomena. Especially, the reduced kinetic energy of the flat bands in twisted bilayer graphene makes it a correlated system. Non-Hermitian strongly interacting Dirac system has recently been calculated [60]. The non-Hermitian physics in quantum many-body systems, such as strongly correlated phases in moié system, will be a future focus.
@@
Note added-. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed [61].
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 12104099 and No. 12274095) and Guangzhou Science and Technology Program (Grant No. 2024A04J0272)

### v3

@@ IV Tripod model and non-Hermitian chiral symmetry
To understand the emergence of non-Hermitian exceptional ring, we use the simplified tripod model [8], which truncates Eq. (1) at the first honeycomb shell. The Hamiltonian is written as
@@
where $h_{tj}$ for $j=1,2$, and $3$ is on the top layer and $h_{b}$ is on the bottom layer. Since $h_{b}$ is non-Hermitian, the Hamiltonian $H_{tri}$ has biorthogonal eigenvectors. In this system, the left and right eigenvectors are four two-component spinors:
@@
Using perturbation theory one can derive the effective Hamiltonian in the space of $\psi_{0}$. The Hamiltonian $H_{tri}$ can be divide into a momentum-independent part $\mathcal{H}^{(0)}$ and a momentum-dependent part $\mathcal{H}^{(1)}_{k}$. The $\psi_{jR}$ can be expressed by
@@
Since $h_{j}$ is independent of non-Hermiticity, we can expect $T_{j}h_{j}^{-1}T_{j}=0$, which is similar to the Hermitian model. Thus, the $\psi_{0}$ spinor in the zero-energy eigenstates satisfies
@@
This form seems to be the same as that of the Hermitian TBG. However, $h_{0}$ is in a non-Hermitian form; the $\psi_{0R}$ is different from $\psi_{0L}$. We can check that the biorthorgonal wave functions are normalized as
@@
The effective Hamiltonian matrix elements to leading order in $k$ is $\langle\Psi^{(i)}|\mathcal{H}^{(1)}_{\textbf{k}}|\Psi^{(j)}\rangle=\psi_{0}^{(i){\dagger}}(-v_{F}^{*}\rm{\sigma}\cdot\textbf{k}+i\lambda_{V}^{*}\sigma_{z})\psi^{(j)}_{0}$, which gives the effective Hamiltonian
$H_{\text{eff}}=-v_{F}^{*}\rm{\sigma}\cdot\mathbf{k}+i\lambda_{V}^{*}\sigma_{z},$ (10)
@@
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to that of the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in the AB region $\alpha_{1}$.
We notice that the effective model in Eq. (10) preserves the non-Hermitian chiral symmetry as
$\mathcal{C}H_{\text{eff}}^{\dagger}(\textbf{k})\mathcal{C}^{-1}=-H_{\text{eff}}(\textbf{k}),$ (13)
where the unitary operator $\mathcal{C}=\sigma_{z}$ satisfies $\mathcal{C}^{2}=1$. This non-Hermitian chiral symmetry is distinct from the Hermitian chiral symmetry due to $H(\textbf{k})\neq H^{\dagger}(\textbf{k})$ [55].
The non-Hermitian chiral symmetry indicates that the eigenvalues come in pairs of ($E,-E^{*}$), which means the two eigenvalues $E_{1}$ and $E_{2}$ of $H_{\text{eff}}(\textbf{k})$ satisfy
$E_{1}=-E^{*}_{2}\text{ or }E_{i}=-E^{*}_{i}\quad for\quad i=1,2.$ (14)
This ensures $\Delta=(E_{1}-E_{2})^{2}$ is real, which is the discriminant of characteristic polynomial $f(E,\textbf{k})=det[E-H_{\text{eff}}(\textbf{k})]$. As we know, if both the real and imaginary parts of $\Delta$ are 00, there is an exceptional point. Since the constraint $Im(\Delta)=0$ is always satisfied in our system, the degrees of freedom (e.g., $k_{x}$ and $k_{y}$) outnumbers the only constraint $Re(\Delta)=0$, which leads to the presence of a higher-dimensional exceptional geometry — the exceptional ring. Thus, the non-Hermitian chiral symmetry protects the exceptional ring in non-Hermitian TBG. A similar mechanism can be found in other two-dimensional non-Hermitian semimetals protected by non-Hermitian chiral symmetry [52, 53, 54, 55].
Figure 3: (color online) At an “exceptional magic angle” $\alpha_{1c}$, the lowest band shows (a) a real part that is extremely flat and (b) an imaginary part with a finite lifetime. The non-Hermitian strength is taken as $\lambda_{V}=0.07$eV. The exceptional magic angle is $\alpha_{1c}=0.46$ away from the Hermitian magic angle $\alpha_{1H}\approx 0.593$.
Figure 4: (color online) (a) The exceptional magic angle $\theta_{c}$ as a mononic increasing function of non-Hermitian strength $\lambda_{V}$ (b) $f(\alpha_{1c})$ as a function of the ratio between the non-Hermitian strength $\lambda_{V}$ and the interlayer hopping $w_{1}$. The red line is the linear fit with a slope of 0.96. The red triangle marks $f(\alpha_{1H})$ in the Hermitian limit.
V Non-Hermitian flat band at exceptional magic angle
In the previous sections, we have shown that an exceptional ring appears at a non-magic twist angle. Here, we show the emergence of a non-Hermitian flat band at an “exceptional magic angle” and discuss the relationship between the exceptional magic angle and non-Hermitian strength considering Fermi velocity renormalization.
For a fixed non-Hermitian strength $\lambda_{V}$, we expect that the exceptional rings cover more regions in the mBZ with decreasing twist angle. This is because the size of the mBZ, $k_{\theta}=2k_{D}\sin(\theta/2)$, decreases with decreasing $\theta$. As the twist angle decreases to a critical value, the three exceptional rings centered at different $K^{\prime}_{M}$s cross each other at $\Gamma_{M}$ and cover the whole mBZ. As shown in Fig. 3, we find that the whole real part of the two lowest bands becomes 00. This means the flat bands are compatible with non-Hermiticity. The presence of these non-Hermitian flat bands in moiré system is the main finding in this work. We can see that these non-Hermitian flat bands are isolated from higher-energy bands, which are distinguished from non-Hermitian flat bands in geometry-frustrated lattices.
The corresponding critical twist angle can be denoted as the “exceptional magic angle”, analogous to the magic angle in the Hermitian case. This exceptional magic angle can be defined by the twist angle at which the bandwidth of the lowest band is minimized. While the bottom of the lowest band at $K^{\prime}_{M}$ from the Dirac point is always at zero energy, the band top should be minimized at the exceptional magic angle. Thus, the band top at the $\Gamma_{M}$ point takes the zero energy value at the exceptional magic angle $\theta_{c}$ and dimensionless $\alpha_{1c}$
$E=\sqrt{(\hbar v^{*}_{F}k_{\theta c})^{2}-\lambda_{V}^{*2}}=0.$ (15)
In Fig. 4(a), we can see that the exceptional magic angle $\theta_{c}$ increases with increasing non-Hermitian strength. This is consistent with our expectation that exceptional rings cover more regions in the mBZ with decreasing the twist angle. Importantly, this offers a feasible way to realize a non-Hermitian flat band at a large twist angle, which is easier to be precisely controlled than at a Hermitian magic angle.
Taking into account the renormalization effect of the Fermi velocity, the relationship between $\alpha_{1c}$ and $\lambda^{*}_{V}$ can be established by using Eqs. (11) and (15). We find that
$\lambda^{*}_{V}=w_{1}f(\alpha_{1c}),\quad\text{with }f(\alpha_{1c})=\frac{1-3\alpha_{1c}^{2}}{(1+3\alpha_{0}^{2}+3\alpha_{1c}^{2})\alpha_{1c}}.$ (16)
In the chiral limit $(\alpha_{0}=0$), $\lambda^{*}_{V}$ can be replaced by $\lambda_{V}$. In Fig. 4(b), we plot $f(\alpha_{1c})$ as a function of $\lambda_{V}/w_{1}$. A linear fit with a slope close to 1 confirms Eq. (16). Also, when $\lambda$ is close to 00, $f(\alpha_{1c})$ approaches $f(\alpha_{1H})$ with a Hermitian magic angle. This is consistent with the Hermitian limit.
Note that when the size of mBZ is large, higher-order momentum terms will give obvious deviation from Eq. (16), since only the leading order in $k$ is considered in Eq. (11). In addition, the band crossing with a higher-energy dispersive band presented at large non-Hermiticity will break the isolation of the non-Hermitian flat band. Both conditions constrain the valid range of non-Hermitian strength.
VI Berry Curvature and Chern number
@@
So far, we have shown the existence of exceptional geometries in the non-Hermitian TBG. In this section, we explore their topological properties numerically.
Figure 5: (color online) Topological charge for the $\sigma_{z}=1$ band in the mBZ as a function of non-Hermitian strength. A small in-plane electric field is imposed to open the gap. The results are in the chiral limit $\alpha_{0}=0$ and the twist angle is taken as $\alpha_{1}=0.384$. While the topological charge remains to be quantized when the non-Hermitian strength exceeds the critical value $\lambda_{Vc}=0.15$ at which the non-Hermitian flat band appears, it fluctuates when the $\sigma_{z}=1$ band crosses with higher bands at large non-Hermitian strength above $\lambda_{V}=0.3$ eV.
The topological stability is ensured by the nontrivial topology for a line gap, when a line can be drawn on the complex energy plane to separate the energies of two bands along a closed path. In our system, the effective Hamiltonian, Eq.(10), preserves the $PT$ symmetry $(PT)H_{\text{eff}}(\textbf{k})(PT)^{-1}=H_{\text{eff}}(\textbf{k})$ with $PT=\sigma_{x}K$. In addition, symmetry-protected exceptional rings with the exceptional geometry dimension $d_{\text{EP}}=1$ in $d=2$ spatial dimension have codimension $p$ defined as $p=d-d_{EP}=1$. From the non-Hermitian topology theory [55], there is a well-defined Chern number corresponding to the $\mathbb{Z}$ index for a line gap for the real part of the energy.
@@
Here, $|\psi^{\beta^{\prime}}(\mathbf{k})\rangle$ and $\langle\psi^{\beta}(\mathbf{k})|$ can be the left or right eigenvectors. Since previous study has proven that the total topological charge is the same for four possible choices of eigenvectors [56], we choose $\beta=L$ and $\beta^{\prime}=R$ in the following calculation.
@@
$\Omega(\mathbf{k})^{\beta,\beta^{\prime}}=\lim_{q\rightarrow 0}\frac{1}{4q^{2}}\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}|\psi^{\beta}_{\mathbf{k}-q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}+q\hat{y}}|\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}+q\hat{x}-q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}-q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}\rangle,$ (19)
@@
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. For small $\lambda_{V}$, we can find that the total Berry charge in one mBZ is 1. This can be easy to understand from two facts. One is that each of the three Dirac points corresponds to Berry curvature monopoles of charge 1 in the absence of non-Hermiticity. The other is that the topological charge is preserved on the exceptional contour that forms from the original Dirac point. Thus, we can expect that each of the three exceptional contours contributes one-third of the Berry charge to one mBZ.
@@
Most interestingly, as shown in Fig. 5, even when $\lambda_{V}$ exceeds the critical non-Hermitian strength $\lambda_{Vc}=w_{1}f(\alpha_{c})$, which corresponds to the exceptional magic angle $\alpha_{c}$, the total Berry charge remains $1$. This is opposite to the merging of exceptional contours with opposite topological charge [36], in which the topological charge is dissipated and gives a single, uncharged exceptional contour. This indicates a breakdown of the Nielsen-Ninomiya theorem and distinguishes the valley topological features of non-Hermitian TBG from a single-layer graphene. This show that the non-Hermitian flat band is topological nontrivial, which distinguishes it from the non-Hermitian flat bands in geometry-frustrated lattices [42, 43, 44, 45, 46]
Note that the band crossings with the remote bands will lead to a fluctuation of topological number as the bands are degenerated. This gives an upper limit of $\lambda_{V}$ for topological non-Hermitian flat bands.
VII Discussion and conclusion
@@
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings are centered at the $K_{M}$ ($K^{\prime}_{M}$) corner in the moiré Brillouin zone in the $K^{\prime}$ ($K$) valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. We found that the exceptional rings are protected by non-Hermitian chiral symmetry. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This non-Hermitian flat band is characterized with a robust Chern number, $\mathcal{C}=1$, in moiré Brillouin zone.
The presence of non-Hermitian flat bands at an “exceptional Dirac magic” angle introduces a new type of band engineering for twisted moiré systems. It is known that the electronic properties of bilayer graphene near a magic twist angle are extraordinarily sensitive to the carrier density and to controllable environmental factors such as the proximity of nearby gates and twist-angle variation [57]. The presence of non-Hermiticity increases the magic twist-angle value and provides some degrees of tuning capability that can reduce the severity of absolute twist-angle control requirements.
@@
The non-Hermitian flat bands in moiré systems are distinguished from those in geometrical frustration lattices, which are embedded in dispersive bands [44]. While the non-Hermitian flat bands in frustrated lattices have been shown to give compact localized states, the wave propagation of non-Hermitian flat bands in moiré system can also be investigated in cold atom and metamaterial systems. Recently, an ultracold atom experiment has had a breakthrough in the realization of atomic Bose-Einstein condensate in twisted bilayer optical lattices  [58]. Also, following the proposals on the photonic analog [59, 60, 61, 62, 63] and the phononic analog [64, 65] of twisted bilayer graphene, demonstration of optical bilayer photonic crystal devices has been recently reported in the microwave range [66] and the optical frequency range [67]. Since on-site dissipation is experimentally realizable in cold-atom and metamaterial systems  [32, 33, 34], the realization of non-Hermitian flat bands in these systems can be anticipated. In particular, the Weyl exceptional ring has been realized in photonic experiments [68, 69].
The non-Hermitian flat band is topological and may lead to the realization of non-Hermitian fractional quantum Hall states. First, non-Hermiticity can enhance strong correlated phenomena. For example, a numerical study on non-Hermitian interacting system on the honeycomb lattice shows that non-Hermitian enhances the antiferromagnetic ordered phase and changes its transition to Dirac semimetal [70]. Second, non-Hermiticity may induce a topological phase that does not have its Hermitian counterpart. For example, the study of non-Hermitian quantum Hall states finds that the best quantization of the non-Hermitian topological invariant is observed at high carrier density and low field [71], which provide efficient operation of devices.
@@
Note added. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed [72].
@@ Acknowledgements.
We thank the helpful discussion with Zhen-Yu Zhang. This work is supported by the National Natural Science Foundation of China (Grants No. 12104099 and No. 12274095) and the Guangzhou Science and Technology Program (Grant No. 2024A04J0272).